# Contextual Paralinguistic Data Creation for Multi-Modal Speech-LLM: Data Condensation and Spoken QA Generation

### v1

@@ 1 Introduction
  • Open-Source Evaluation Dataset: We provide an open-source benchmark comprising 480 speech samples paired with CPQA pairs. [^1] The dataset link.
@@ 3.1 Preliminary study for framework parameters
We apply length filtering with a minimum duration of $\tau=30$ sec. For sub-segment SER consistency filtering, we explore UWA across different valence thresholds. Unlike accuracy, UWA mitigates selection bias toward dominant categories, addressing the varying difficulty of emotion categories in SER. Thresholds in $X_{E}$ are set as $v_{\text{pos,min}}=x,v_{\text{neg, max}}=1-x,v_{\text{neu, min}}=y$, and $v_{\text{neu,max}}=1-y$. Figure 3 shows the highest UWA 33.65% is achieved at $x=0.5$ and $y=0.4$ on the SG TV/movie dataset. For occurrence filtering, we set $\alpha=[10,10,4,4,2,3]$ for the six non-neutral emotion categories: angry, disgusted, fearful, happy, sad, and surprised, to ensure retain sufficient and balanced samples for each category in the condensed dataset.

### v2

@@ 1 Introduction
  • Open-Source Evaluation Dataset: We provide an open-source benchmark comprising 480 speech samples paired with CPQA pairs. [^1] https://huggingface.co/datasets/MERaLiON/CPQA-Evaluation-Set
@@ 3.1 Preliminary study for framework parameters
We apply length filtering with a minimum duration of $\tau=30$ sec. For sub-segment SER consistency filtering, we explore UWA across different valence thresholds. Unlike accuracy, UWA mitigates selection bias toward dominant categories, addressing the varying difficulty of emotion categories in SER. Thresholds in $X_{E}$ are set as $v_{\text{pos,min}}=x,v_{\text{neg, max}}=1-x,v_{\text{neu, min}}=y$, and $v_{\text{neu,max}}=1-y$. Figure 3 shows the highest UWA 33.65% is achieved at $x=0.5$ and $y=0.4$ on the SG TV/movie dataset. For occurrence filtering, we set $\alpha=[10,10,4,4,2,3]$ for the six non-neutral emotion categories: angry, disgusted, fearful, happy, sad, and surprised, to ensure sufficient and balanced samples are retained for each category in the condensed dataset.
@@ 5 Summary
6 Acknowledgement
This research/project is supported by the National Research Foundation, Singapore, under its National Large Language Models Funding Initiative. Any opinions, findings, conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore.